X+2 1
--- = - -[X+2]
3 3

Answer by checkley71(8403)

X+2 1
--- = - -[X+2] cross multiply
3 3
3(x+2)=3*-(x+2)
3x+6=-3x-6
3x=3x=-6-6
6x=-12
x=-12/6
x=-2 answer.
proof:
(-2+2)/3=-(-2+2)/3
0/3=0/3
0=0